Burt is 14 years older that Erney. In eight years, twice Erneys age will be six more than what Burt's age will be. Find the age of Burt now.
Answer by ionesco(12)

we have 2 equations:
B=14+E
2(E+8)=6+(B+8)
then: Erney's age will be 12 and Burt's 26